jimmy is 36 years old and james is 11 years old.In how many years, will jimmy's age be twise the age of james?
Answer by macston(5194)

.
x=number of years until Jimmy is twice as old as James.
.
36+x=2(11+x)
36+x=22+2x
14=x ANSWER: In 14 years Jimmy will be twice as old as James.
.
CHECK:
Jimmy is 36, in 14 yrs will be 50 years old.
James is 11, in 14 yrs will be 25 years old.
In 14 years, 50 years is twice 25 years.